OpenStudy (anonymous):

can anyone help me in finding the area between two curves, I got wrong answer , (68/1+x^2) and y=|x|

OpenStudy (anonymous):

34pi-4 is my answer

OpenStudy (mathmate):

Very close, probably some minor mistake in the integration. Can you post your integrand and limits of integration?

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \tan^{-1} (x)-x^2\]

OpenStudy (mathmate):

I would like to see the integrand, i.e. before integration or any substitution. Also, did you make a sketch of the functions, and determined the intersection points?

OpenStudy (anonymous):

yes , 1 and -1

OpenStudy (mathmate):

Well, I get limits of -4 and 4, which also shows by inspection that both functions are even. This means that we can integrate from 0 to 4 and double the value for the answer.

OpenStudy (mathmate):

Apart from the limits, you may consider the integrand as 68/(1+x^2)-x. What you gave me could be some part of the antiderivative.

OpenStudy (anonymous):

so instead of a=0 b=1 , I must use a=0 b=4?

OpenStudy (mathmate):

yes, something like: \(2\int_0^4(\frac{68}{1+x^2}-x)dx\)

OpenStudy (anonymous):

thanks @mathmate , last question , how did you get the -4 and 4 ? when I am solving it I always ger 1, -1 as points of intersection

OpenStudy (anonymous):

thanks , I already got the answer , I clicked wrong number , thanks anyway

OpenStudy (mathmate):

ok, you're welcome!

OpenStudy (anonymous):

so the correct answer is 164.3112023? 136(tan^-1(4)-tan^-1(0))-1?

OpenStudy (anonymous):

136(tan^-1(4)-tan^-1(0))-4^2? am I right @mathmate ?

OpenStudy (mathmate):

Use \(tan^{-1}0 = 0\) to simplify your answer. Also the integral of x is \(\frac{x^2}{2}\). Do not forget to multiply by two for the final answer (136 is correct).